### v2

Abstract
Federated graph learning is an emerging field with significant practical challenges. While many algorithms have been proposed to enhance the accuracy of training graph neural networks, e.g., for node classification problems on large graphs, in a federated manner, their system performance is often overlooked, even though it is crucial for real-world deployment. To address this gap, we introduce FedGraph, a research library built for practical distributed deployment and benchmarking in federated graph learning. FedGraph supports a range of state-of-the-art graph learning methods and includes built-in profiling tools to evaluate system performance, focusing specifically on communication and computation costs during training. Unlike existing benchmark platforms, FedGraph natively incorporates homomorphic encryption to enhance privacy preservation and facilitates the development of practical applications by enabling distributed training across multiple physical machines, providing an evaluation framework that can guide the system design of future federated graph learning algorithms. Leveraging these optimizations, we use FedGraph to demonstrate the first privacy-preserving federated learning system to run on graphs with 100 million nodes.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
Graph neural networks aim to learn representations of graph-structured data that capture features associated with graph nodes as well as edges between the nodes Bronstein et al. (2017). Such representations can be used to represent and classify nodes or predict the formation of links in large graphs, or can be used to represent and classify smaller graphs. Popular applications of these learning tasks range from recommendation systems (e.g., predicting link formation in graphs if links represent user preferences for products) and fake news detection in social networks (e.g., classifying nodes in a social network graph as “fake” where nodes represent news items) to drug discovery in biology (e.g., representing cells as graphs and classifying them according to drug susceptibility) and fraud detection in sensor networks (e.g., classifying nodes as “fraudulent” if the nodes represent sensors connected together in a graph) Benamira et al. (2019); Zhang et al. (2020b).
Real-world graph data is often too large to be trained on a single server or may naturally exist on multiple local clients. For example, graph learning on records of billions of users’ website visits requires significant computational resources beyond those of a single server. Even if a single server or data center could hold such information, privacy regulations may require that it be stored in the computers or smartphones at which it was generated. Strict data protection regulations such as the General Data Protection Regulation (GDPR) in Europe and Payment Aggregators and Payment Gateways (PAPG) in India, for example, prevent private user data from being shared across international borders. Users may also not want to share their personal data with the server. Federated learning then shows promise for preserving user privacy while training accurate models on data stored at multiple clients Zhao et al. (2018).
These challenges to training on real-world graphs motivate federated graph learning (FGL) as an important research topic Yao et al. (2024); Xie et al. (2021); Zhang et al. (2021). The training schematic of such a federated architecture is shown in Figure 1: each client maintains a local graph and a local graph neural network (GNN) model, which can be aggregated with other clients’ local models at a coordinator server. Within this architecture, there are many different FGL algorithms that handle local model updates and aggregation differently, resulting in varying model accuracies and system performance (e.g., runtime, communication costs) Yao et al. (2024); Xie et al. (2021); Zhang et al. (2021). While evaluating the model accuracy of such algorithms can be easily done with simulations on open graph datasets Hu et al. (2020), evaluating their real-world system performance requires sophisticated benchmarking platforms that can fairly compare multiple FGL algorithms.
Vanilla FGL benchmark platforms are designed for research purposes only and do not have a distributed training system, instead of simulating the actions of multiple physically distinct clients on a single machine. Thus, they do not capture the real communication and computation costs of training graph models in a federated manner. Meanwhile, existing federated learning libraries mainly focus on overcoming client data heterogeneity with limited support for GNN training. For example, FedScale Lai et al. (2022) does not support graph models, while FedGraphNN He et al. (2021) and FederatedScope-GNN Wang et al. (2022) are not well maintained and have limited support of multiple FGL algorithms, forcing users to implement the FGL algorithms that they desire to compare on these benchmark platforms themselves. Moreover, none of these platforms natively support FGL enhancements that may be needed in practice, e.g., homomorphic encryption for enhanced privacy preservation Jin et al. (2023). Table 1 compares their features to our FedGraph library.
Figure 1: Federated Graph Training Schematic. Nodes in a graph (shown as circles) are distributed across clients. Arrows indicate that each client can exchange model updates with a central server during the training process of a graph neural network model.
Vanilla FGL FedGraphNN FedScale FederatedScope-GNN FedGraph
Distributed Training
Graph Learning
Multiple FGL Algorithms
Cross-Client Edges
Encrypted Aggregation
System Level Profilor
Large Graph Node Classification
Table 1: FedGraph Comparison with Former Frameworks. FedGraph supports distributed training with various FGL algorithms. It then supports algorithms for overcoming cross-client edges and encrypted aggregation. With a system-level profiler, FedGraph can then benchmark existing algorithms and provide optimizations to support node classification on large graphs.
To meet these shortcomings, we introduce FedGraph (Federated Graph) [^1] https://github.com/fedgraph/fedgraph, a research library to easily train GNNs in federated settings. As shown in Table 1, FedGraph supports various federated training methods of graph neural networks under both simulated and real federated environments, as well as encrypted communication between clients and the central server for model update and information aggregation. We summarize our contributions in FedGraph as follows.
  • We introduce FedGraph, a comprehensive Python library built for real-world federated graph learning systems, incorporating multiple state-of-the-art FGL algorithms to facilitate easy comparison between them.
  • FedGraph provides an advanced system-level profiling tool that enables detailed analysis of communication and computation costs throughout training, as well as encrypted aggregation to strengthen privacy protection.
  • Experimental benchmarks highlight FedGraph’s effectiveness in addressing key challenges in federated graph learning. We demonstrate that FedGraph enables detailed comparisons of different FGL algorithms under various configurations and showcases its ability to support a privacy-preserving federated system on graphs with 100 million nodes.
We give an overview of FedGraph’s system design in Section 2 before detailing our implementation of FedGraph’s components in Section 3. We discuss the usability of FedGraph as an evaluation platform in Section 4 and present examples of the experiments that it enables in Section 5. We then conclude in Section 6.
2 FedGraph System Design
In this section, we first introduce the three main graph learning problems that FedGraph supports and highlight the design principles of the library and how we achieve them.
2.1 Graph Learning Problems
Most FGL applications can modeled as one of three major graph learning problems: node classification (e.g., classifying nodes representing papers in citation networks based on their research topic), link prediction (e.g., recommending the formation of links that represent friendship between users in social networks), and graph classification (e.g. classifying types of proteins in biology, where each protein is represented as a graph). Figure 2 illustrates the three major aspects as different graph tasks.
Figure 2: Modeling Applications as Graph Tasks. Nodes are connected by edges, where the color of the node represents the information of the node. Node classification (left) is to classify the type of grey nodes based on the graph structure and the information from other nodes. Link prediction (middle) is to predict the future links between nodes based on existing edge and node information. Graph classification (right) is to classify the type of the graph based on the graph structure.
Given these three learning problems, FedGraph supports running algorithms that aim to solve the following learning tasks:
Federated Node Classification with Cross-Client Edges: A key challenge in node classification is the presence of cross-client edges that connect nodes at different clients Yao et al. (2024). Thus, the library supports communicating information about these nodes stored in other clients, either before or during training, without affecting the privacy of users.
Federated Link Prediction on Dynamic Graphs: A key challenge in link prediction is that links may form at clients at different rates. Thus, the library supports balancing temporal heterogeneity across clients with privacy preservation.
Federated Graph Classification on Heterogeneous Data: The library supports various algorithms for graph classification with non-IID distributions of graphs across clients.
2.2 Design Principles
The library is designed for a federated learning researcher or applied scientist, as shown in the design diagram of FedGraph in Figure 3. Our design follows four principles:
Figure 3: Design Diagram of FedGraph. The system is organized into four layers: the user access layer, the application layer, the domain layer, and the infrastructure layer. Users only need to focus on the access layer while the developers can focus on one of the remaining layers based on the domain knowledge.
Easy-to-use and unified API: In the access layer, users first set up the training configuration and then use the same code for local simulation and federated training. All it takes is 10-20 lines of code to get started with training a federated GNN model with the same code for both simulation and distributed training. GNN models are PyTorch models provided by PyG and DGL. The federated training process is handled by Ray. We abstract away the complexity of federated graph training and provide a unified API for training and evaluating FedGraph models.
Various federated graph learning methods: In the application layer, the library supports three tasks with various algorithms and datasets. Most of the state-of-the-art federated graph training methods have been implemented by library developers or authors of research papers and are ready to be applied.
Great flexibility on new datasets and new algorithms: For the domain layer, we separate the code into data loader and federated training classes for better management. Existing FedGraph models can easily be extended for conducting your research. Simply inherit the base class of trainers and implement your methods.
Large-scale real-world federated graph training: In the infrastructure layer, by utilizing Ray and Kubernetes, the library can execute processes and store data in various edge/cloud computation devices. Homomorphic encrypted computation can be enabled to further enhance the federated aggregation. We can then meet the demand for FedGraph applications in challenging real-world scenarios with privacy preservation and support learning on large-scale graphs across multiple clients.
3 FedGraph System Components
In this section, we first present the FedGraph API for running different federated graph learning tasks. We then introduce FedGraph Homomorphic Encryption for privacy-preserved aggregation and FedGraph Kubernetes for scalable distributed training.
3.1 FedGraph API
When calling the FedGraph API, as shown in Figure 4, Users can simply specify the name of the dataset and algorithm, the API will then call the corresponding data loader class to generate the required data, which is then automatically fed into the appropriate algorithm runner for tasks like Node Classification, Graph Classification, or Link Prediction. Additionally, users can seamlessly add their own datasets or federated graph learning algorithms if needed, as long as they satisfy the form requirements for the specified task.
Figure 4: High-level Demonstration of FedGraph API Design
3.2 FedGraph Homomorphic Encryption
FedGraph supports homomorphic encryption (HE) to ensure end-to-end secure computation to preserve user privacy across various graph learning algorithms, as shown in Figure 5. To our knowledge, this is the first federated graph learning library to incorporate end-to-end homomorphic encryption. While previous work explored homomorphic encryption in specific federated graph learning context Yao et al. (2024), FedGraph advances this approach through several key innovations that address challenges:
1. The system implements HE in different stages of pre-training and training based on algorithm design, enabling flexible security protocols across different FGL tasks.
2. The system maintains an efficient implementation of the homomorphic encryption process.
3. The system ensures secure aggregation for both feature and structural information in graph data
Figure 5: Homomorphic Encryption Process. The clients own the private key and the server is used to perform secure aggregation without knowing the plaintext data.
Pre-Train Aggregation: During pre-training, each client maintains its local node features and needs to share the aggregated neighbor information without revealing the raw data. Such information is encrypted before sending it to the server. The server then performs encrypted aggregation and then sends the result back to clients. The client then decrypts the feature aggregation without knowing the original feature.
Federated Aggregation During Training: During model aggregation on the server side, HE can also ensure secure aggregation during federated training without revealing fine-grained local updates to the server Zhang et al. (2020a); Jin et al. (2023). For the training process, we encrypt gradient updates before transmission for training parameters. Then, the central server performs aggregation on the encrypted values and homomorphic averaging on the aggregation. Such HE aggregation provides robust protection while maintaining computational efficiency.
3.3 FedGraph Kubernetes
We employ Kubernetes to optimize the scalability and reliability of the application layer, allowing our federated learning system to handle dynamic workloads effectively. As illustrated in Figure 6, our approach utilizes Amazon Web Services (AWS) Elastic Kubernetes Service (EKS) to facilitate flexible resource management, deploying a self-managed Kubernetes cluster equipped with GPU instances to support computationally intensive tasks. The architecture comprises a master node that oversees orchestration responsibilities, including task scheduling, resource allocation, and monitoring of cluster health, while worker nodes are responsible for executing distributed processes in parallel. To further enhance scalability and efficiency, we integrate Kubernetes Cluster Autoscaler, which automatically adjusts the number of worker nodes in response to fluctuating workload demands. This setup ensures that resources are provisioned only when necessary, leading to optimized costs and performance. The result is a robust, scalable federated learning environment that maintains high efficiency even under varying levels of load, supporting continuous, large-scale learning processes.
Figure 6: Kubernetes Service for FedGraph Computation Cluster Management
3.4 FedGraph Monitoring System
The FedGraph Monitoring System, represented in Figure 7, uses a Monitor Class to track CPU, GPU usage, memory, and communication costs between server and trainer nodes. Here is an overview of the monitoring system:
CPU/GPU Usage Monitoring: The Monitor Class keeps track of resource consumption within the Server Process and Ray Pod AutoScaler, providing insights into processing efficiency and identifying when additional resources might be required.
Memory Usage Tracking Memory usage across the API Gateway, Object Store, and server processes are monitored to detect potential bottlenecks, helping to ensure smooth operation and efficient memory allocation.
Communication Cost Logging: Data transfer rates between the FedGraph User and API Gateway are recorded to measure network performance and detect any latency issues affecting responsiveness.
This system optimizes resource use, tracks performance trends, and helps identify bottlenecks in real time.
Figure 7: FedGraph Monitoring System Architecture
3.5 Scalability
The FedGraph system is built to handle large-scale federated learning, supporting thousands to millions of nodes through the following key design features:
Client Selection for Large-Scale Federated Learning: FedGraph uses a selective client engagement mechanism, only partial clients send the model update at each training round. This approach reduces communication costs and resource strain, enabling scalability to thousands of clients per training round.
Minibatch Training for Federated Updates: Minibatch training allows each client to process only a subset of the local graph, reducing computational and memory demands. This enables devices with limited resources to participate, improving convergence speed without overwhelming network resources.
Communication and Resource Optimization: FedGraph optimizes communication, minimizing data transfer between clients and the server. Additionally, Kubernetes-based dynamically manages Server Pods and Trainer Pods to meet workload demands, ensuring efficient resource usage as the system scales.
These features collectively enable FedGraph to support federated learning at a massive scale while maintaining resource efficiency and performance.
4 FedGraph Usability
In this section, we first an example of how to use the library, then show the configuration of homomorphic encryption and measurement dashboard.
4.1 FedGraph Use Example
Researchers and applied scientists can easily install the library, set up configurations, and then start training.
{python}
## Installation pip install fedgraph
After installing the library, the user can run the experiment with the following steps:
Task specification: The user declares the task on which the runner will focus. As we have involved multiple graph learning algorithms that cannot share the same configuration, we require the user to further specify the algorithm they want to deploy instead of manually modifying the configuration file.
Configuration loading: The configuration file contains all the necessary information needed for the training, including dataset details, model hyperparameters, and some general training options. The configuration is set up for different tasks and algorithms.
Data loading: Once started running, a build-in module data_loader is used for loading the data if a correct set of configurations is supplied. We also provide the data interface to support customized datasets.
Distributed graph training: Once the data gets loaded into different clients, the training experiment starts by calling the general runner run_fedgraph.
The implementation details of the API usage are demonstrated in the following code block.
{python}
from fedgraph import data_process from fedgraph import federated_methods import attridict
config = ”dataset”: ”cora”, ”fedgraph_task”: ”NC”, ”method”: ”FedGCN”, ”global_rounds”: 100, ”local_step”: 3, ”learning_rate”: 0.5, ”n_trainer”: 2, ”num_layers”: 2, ”num_hops”: 2, ”iid_beta”: 10000, ”gpu”: False, ”encrypted_computation”: True, ”logdir”: ”./runs”,
config = attridict(config) data = data_process.data_loader(config) federated_methods.run_fedgraph(config,data)
4.2 Homomorphic Encryption Configuration
We provide a comprehensive guide for the configuration of Homomorphic Encryption in FedGraph to provide security guarantees corresponding to the graph structure and size.
4.2.1 Parameter Configuration
In Table 2, we present the key parameters for configuring CKKS homomorphic encryption. The selection of the parameters is based on the dataset size and desired security level. Different parameter combinations create tradeoffs between computational overhead, communication cost, and precision. We present the default setting and parameters selection range to guide the user in selecting an appropriate combination that achieves the balance between objectives.
Parameter Default Value Description Range
scheme CKKS Encryption scheme type N/A
polynomial modulus degree 16384 Maximum degree of polynomials used to represent encrypted data ($N\geq 2\times$ max(nodes, features)) 4096, 8192, 16384, 32768
coefficient modulus bit size [60, 40, 40, 40, 60] Bit size for coefficient modulus that controls precision Array of integers in the range $[20,60]$
global scale $2^{40}$ Global scale factor for encoding precision $2^{30}$, $2^{40}$, $2^{50}$, etc.
security level 128 Bit security level 128, 192, 256
Table 2: TenSEAL Homomorphic Encryption Configuration Parameters. This table shows the key parameters for configuring the CKKS encryption scheme in the FedGraph library, including their default values, descriptions, and available ranges.
Method Poly _mod Coeff_mod Precision Dataset Time(s) Comm_cost (MB) Accuracy
FedGCN (plaintext) N/A N/A N/A Cora 13.29 59.21 0.783 $\pm$ 0.07
FedGCN (HE) 16384 [60,40,40,40,60] $2^{40}$ Cora 27.71/23.22/56.05 3279.15 0.779 $\pm$ 0.08
FedGCN (HE) 32768 [60,40,40,40,60] $2^{50}$ Cora 29.44/36.17/71.76 4434.58 0.781 $\pm$ 0.08
Table 3: Microbenchmark of FedGCN under Homomorphic Encryption with different CKKS scheme parameters.
4.2.2 Microbenchmark
We then provide the microbenchmark of HE on federated graph training in Table 3. The experiments are conducted on a 2-layer FedGCN for node classification tasks, running 100 global rounds with default settings in Cora. For CKKS parameters, we evaluate different polynomial modulus degrees (Poly_mod), coefficient modulus sizes (Coeff_mod), and precision levels. Time(s) for the encrypted version shows pre-train/training/total times respectively. Communication costs (Comm_cost) include both pre-training and training rounds. The complete version can be found in Table 6 in Appendix.
Dynamic Precision: We adjust encryption parameters based on graph sizes and the numerical precision needed. For graphs like Cora, a polynomial modulus degree of 16384 with precision $2^{40}$ satisfies the ideal accuracy, while increased value provides more precise security protection.
Communication Cost Optimization: We employ several strategies to manage the communication overhead inherent in HE operation. The selection of coefficient modulus chain, [60,40,40,40,60], etc., enables efficient depth management for multiple HE operations. Depending on specific dataset characteristics (sparse matrix, larger datasets, etc.), we also employ efficient encryption methods to optimize communication cost and balance the performance.
4.3 System Measurement Dashboard
In Figure 8, the FedGraph project monitors system performance by integrating Prometheus and Grafana within a Kubernetes cluster. Prometheus collects metrics from Ray clusters, while Grafana visualizes these metrics through embedded panels, allowing real-time monitoring and analysis.
Figure 8: FedGraph Monitoring System.
5 System Benchmark Results
In this section, we provide the system benchmark for graph classification, node classification, link prediction, and system scalability.
Figure 9: Accuracy, Communication Cost and Train time of Different Graph Classification Algorithms (10 clients)
Figure 10: Pretrain, Train time, Communication Cost, Accuracy of FedGCN on Ogbn-Arxiv in Different Batch Sizes (200 rounds)
5.1 Graph Classification
The graph classification experiment aims to evaluate the effectiveness of various federated learning algorithms in classifying graph data across diverse datasets, including IMDB-BINARY, IMDB-MULTI, MUTAG, BZR, and COX2. Each dataset represents different graph structures, such as molecular and social networks, providing a comprehensive testbed for algorithm performance.
In Figure 9, our evaluation process spans 200 rounds, accuracy results indicate that GCFL+ and GCFL+dWs consistently achieve higher accuracy across datasets, with MUTAG and COX2 showing the best results across all algorithms. FedAvg and GCFL exhibit moderate performance, with accuracy varying based on the dataset’s complexity. GCFL+ and GCFL+dWs have relatively high training times, particularly on complex datasets like MUTAG and BZR. FedAvg shows the shortest training time across datasets, making it more suitable for time-constrained applications.
As shown in the right panel, GCFL incurs the highest communication cost, particularly for BZR, which highlights a trade-off between communication efficiency and algorithmic complexity. This suggests that while GCFL may offer performance advantages, it does so at the expense of higher data exchange. In contrast, FedAvg demonstrates the lowest communication cost, making it a more suitable choice for scenarios with network constraints where minimizing communication is essential.
5.2 Node Classification
In our node classification experiments, we evaluate the impact of different label distribution metrics using the IID Beta setting. We utilize varying values of $\beta$ (10, 100, 10000) to control the data distribution, where larger $\beta$ values represent more IID (independent and identically distributed) scenarios, and smaller values indicate increasingly non-IID distributions. This approach allows us to simulate diverse federated environments and study the algorithms’ performance under varying data conditions.
Our experiments span multiple datasets like Cora, Citeseer, ogbn-arxiv. With ogbn-products and ogbn-papers100M serving as large-scale benchmarks, helping to further understand model scalability.
We also assess several node classification algorithms, including FedAvg and FedGCN, across different batch sizes. This analysis provides insight into the algorithms’ adaptability and efficiency with both IID and non-IID data.
In terms of system monitoring, we tracked key performance metrics across 10 training nodes and a server node. These metrics include test accuracy, training time, memory consumption, and network usage across both pretrain and train phases.
Figure 11: Training Time, Communication Cost, Test Accuracy on Ogbn-Arxiv in Large Number of Clients
Figure 12: Training Time, Communication Cost, Test Accuracy on Ogbn-Papers100M in Different Batch Sizes (800 rounds)
Figure 13: Comparison of FedGCN Execution Time (top) and Communication Cost (bottom) with HE across Different Datasets
Figure 14: Training time, Communication Cost, AUC Comparison of Different LP Algorithms
5.3 Mini-Batch for Large Dataset
FedGraph inherently supports a mini-batch graph splitter, enabling efficient graph-level processing. Our experiments on the ogbn-arxiv dataset examined the impact of varying batch sizes (16, 32, 64) on test accuracy, training time, memory usage, and communication cost across 10 clients in a federated setting.
Results in Figure 10 show test accuracy peaks at a batch size of 64, with improvements leveling off beyond 32. Training time and memory demands increase with batch size, though pretrain memory remains stable.
5.4 Scalability: Increase Number of Clients
To better test scalability and fit real-world data, we increase the number of clients to 1000. In Figure 11, we observe that as the number of clients increases, the overall training time grows significantly due to added communication overhead and the need for increased synchronization among clients.
As we scale from 10 to 1000 clients under a fixed IID Beta value, there is a small decline in accuracy, likely due to the increased data heterogeneity each client possesses. The communication cost also escalates notably with more clients, highlighting the trade-off between parallelism and efficiency in federated settings. This experiment underscores the system’s ability to handle large-scale client distributions while revealing the resources required to maintain accuracy and efficiency.
5.5 Scalability: Real-World Dataset with the 200 countries population
In our experiments on the ogbn-papers100M dataset over 50GB, we used Hugging Face to store and split the dataset, utilizing the Trainer class to manage local loading for each client. We distributed the data across 195 clients, adjusting the node count for each client to fit a power-law distribution based on country population sizes. This approach mimics real-world federated environments, where larger clients receive more data, accurately reflecting varying data availability. Figure 12 illustrates three key metrics as we vary the batch size (16, 32, and 64).
While training times are consistent across batch sizes, pretrain times increase proportionally with batch size, indicating a need for efficient preprocessing management.
Memory usage remains stable (about 17.5GB), showing the model’s memory footprint is well-optimized. Test accuracy is largely unaffected by batch size, with a slight increase from 16 to 32 that levels off at 64, suggesting flexibility in batch size selection without compromising accuracy.
5.6 Homomorphic Encryption
We conduct microbenchmarks through different selections of parameters and methods to evaluate the performance metrics and trade-offs of HE on FedGCN in Table 3. When comparing among datasets, we observe that HE maintains equivalent accuracy across different parameter selections, as long as they satisfy the modulus requirement. If a smaller than required parameter size is used, the accuracy drops sharply which indicates not valid encryption.
We present Figure 13 to visualize the overhead that HE adds to FedGCN. It illustrates the execution time comparison between plain-text and encrypted versions across different datasets. The computational overhead increases particularly with larger datasets. When comparing FedGCN running across different datasets, feature aggregation takes up a larger portion of communication cost with increasing dataset size. This could be due to the difference in feature and model parameter aggregation in FedGCN, as larger datasets usually have more complex feature representations that require more extensive encryption overhead.
5.7 Link Prediction
In the federated link prediction setting, each client holds region-specific data, respecting privacy constraints by avoiding raw data sharing. The algorithms are assessed on AUC for accuracy, training time for computational efficiency, and communication cost for network efficiency in a distributed, region-specific setup. In Figure 14, the experiment shows that STFL and FedLink achieve the highest AUC values across datasets, with StaticGNN and 4D-FED-GNN+ performing moderately well, but notably lower for the simplest dataset (US). For train time, FedLink and STFL exhibit the longest train times, especially on the largest dataset. 4D-FED-GNN+ has the shortest training times, showing efficiency for quicker iterations. FedLink has the highest communication cost, particularly on the largest dataset, while StaticGNN+ maintains the lowest cost across all datasets, making it more network-efficient.
6 Conclusion
In this paper, we introduce FedGraph, a Python library designed for benchmarking federated graph learning algorithms. Unlike prior evaluation platforms, FedGraph is designed to specifically support multiple FGL algorithms, facilitating their comparisons under different system, algorithm, and dataset configurations. FedGraph supports truly distributed training and homomorphic encryption, allowing for real-world evaluation of communication and computing costs for private FGL. The FedGraph API and system profiler facilitate easy use by machine learning researchers, and the library also allows users to easily introduce their own datasets and algorithms for evaluation. We showcase FedGraph’s ability to support FGL algorithm comparisons and private FGL on large graphs with 100 million nodes, demonstrating that it enables real-world FGL evaluation.
Appendix A FedGraph Code Structure
Fedgraph library can be separated into six modules:
Data Process Module: data_process.py. This module is responsible for data generation and processing. It should be called before calling the runner, so it is only applicable for node classification and graph classification in the latest version. (In link prediction, the dataset is generated and processed inside the runner.)
Runner Module: federated_methods.py. Task-based runners are defined in this module. For node classification and link prediction, there’s a shared runner that could call inner modules to perform federated graph learning. For graph classification, it will further assign the program to an algorithm-based runner.
Server Classes Module: server_class.py. It defines different task-based server classes including server_NC, server_GC, and server_LP.
Trainer Classes Module: trainer_class.py. It defines different task-based trainer(client) classes including trainer_NC, trainer_GC, and trainer_LP.
Backbone Models Module: gnn_models.py. It defines different backbone model classes. Generally, it is task-based (i.e., each task corresponds to one backbone model), but for some models, their variants are also included and the user could also switch the backbone model or define a new one by itself.
Utility Functions: utils_nc.py, utils_gc.py, utils_lp.py. In the current version, the utility functions for different tasks are located in separate Python modules, which is convenient for developing. In the later versions it might be better to use the shared ‘utils.py‘, and use different markers like “NC”, “GC”, and “LP” to distinguish them.
Appendix B Run_fedgraph
The function run_fedgraph is a general runner that receives the dataset and the configurations. It will further assign the program to a task-specified runner run_NC , run_GC , or run_LP based on the user’s specification on Task.
{python}
def run_fedgraph(args, data): if args.fedgraph_task == ”NC”: run_NC(args, data) elif args.fedgraph_task == ”GC”: run_GC(args, data) elif args.fedgraph_task == ”LP”: run_LP(args)
Appendix C Supported Algorithms and Datasets
For federated node classification, link prediction, and graph classification, we integrated different datasets and algorithms for each task, shown in Table 4 and Table 5. Researchers can also easily implement new algorithms and add their datasets.
Task Dataset
Node Classification
Cora, Citeseer, Pubmed,
Ogbn-Arxiv, Ogbn-Products, Ogbn-MAG Hu et al. (2020)
Graph Classification
MUTAG, BZR, COX2, DHFR,
PTC-MR, AIDS, NCI1, ENZYMES,
DD, PROTEINS, COLLAB,
IMDB-BINARY, IMDB-MULTI Xie et al. (2021)
Link Prediction FourSquare Yang et al. (2019), WyzeRule Kamani et al. (2024)
Table 4: Supported datasets of node classification, graph classification, and link prediction in federated learning.
Task Algorithm Backbone
Node Classification FedAvg Li et al. (2019) GCN
Distributed GCN GCN
BNS-GCN Wan et al. (2022) GCN
FedSage+ Zhang et al. (2021) GraphSage
FedGCN Yao et al. (2024) GCN,GraphSage
Graph Classification SelfTrain GIN
FedAvg Li et al. (2019) GIN
FedProx Li et al. (2020) GIN
GCFL Xie et al. (2021) GIN
GCFL+ Xie et al. (2021) GIN
GCFL+dWs Xie et al. (2021) GIN
Link Prediction FedAvg Li et al. (2019) GCN
STFL Lou et al. (2021) GCN
FedGNN+ Gürler & Rekik (2022) GCN
FedLink GCN
FedRule Yao et al. (2023) GCN
Table 5: Supported algorithms of node classification, graph classification, and link prediction in federated learning.
Appendix D Runner Workflow
D.1 Runner Workflow for Graph Classification Task
In the graph classification task, the dataset name can be specified by the user, and it could be either a single dataset or multiple datasets. For single dataset GC, the graphs will be assigned to a designated number of clients; for multiple datasets GC, each dataset will correspond to a client. Dataset generation and preparation are controlled by the function dataloader_GC in the module data_process.py. It will further assign the data process task to the function data_loader_GC_single or data_loader_GC_multiple. All the provided datasets are built-in TUDatasets in the Python library torch_geometric.
[…]
D.2 Runner Workflow for Link Prediction Task
For the link prediction task, we provide a common dataset. The user only needs to specify the country codes. The original datasets are stored in Google Drive, so the user does not need to prepare the dataset by itself. The API will automatically check whether the dataset already exists and download the corresponding one if not.
There are also multiple available algorithms for the link prediction task. However, they share the same set of arguments so that we don’t need to create separate .yaml files. The user could conveniently select different algorithms by directly changing the algorithm field in the configuration file. All the algorithms will share the same runner. The whole workflow is demonstrated in Figure 15(c).
(a) Node classification.
(b) Graph classification.
(c) Link prediction.
Figure 15: Federated graph training workflows for (a) node classification, (b) graph classification, and (c) link prediction tasks.
Appendix E Microbenchmark for FedGCN
We present the full version of Microbenchmark for FedGCN in Table 6.
Method Poly _mod Coeff_mod Precision Dataset Time(s) Comm_cost (MB) Accuracy
FedGCN (plaintext) N/A N/A N/A Cora 13.29 59.21 0.783 $\pm$ 0.07
FedGCN (HE) 16384 [60,40,40,40,60] $2^{40}$ Cora 27.71/23.22/56.05 3279.15 0.779 $\pm$ 0.08
FedGCN (HE) 32768 [60,40,40,40,60] $2^{50}$ Cora 29.44/36.17/71.76 4434.58 0.781 $\pm$ 0.08
FedGCN (plaintext) N/A N/A N/A Citeseer 20.39 187.99 0.658 $\pm$ 0.06
FedGCN (HE) 8192 [60,40,40,60] $2^{40}$ Citeseer 79.35/30.63/113.08 5791.42 0.660 $\pm$ 0.07
FedGCN (HE) 16384 [60,40,40,40,60] $2^{40}$ Citeseer 123.22/45.8/173.40 8084.50 0.652 $\pm$ 0.06
FedGCN (plaintext) N/A N/A N/A PubMed 25.78 150.43 0.774$\pm$ 0.12
FedGCN (HE) 8192 [60,40,40,60] $2^{40}$ PubMed 70.28/19.73/93.18 3612.60 0.757 $\pm$ 0.19
FedGCN (HE) 16384 [60,40,40,40,60] $2^{40}$ PubMed 123.22/45.8/173.40 8084.50 0.769 $\pm$ 0.13
Table 6: Microbenchmark of FedGCN under Homomorphic Encryption with different CKKS scheme parameters. The experiments are conducted on a 2-layer FedGCN for node classification tasks, running 100 global rounds with default settings across three datasets (Cora, Citeseer, PubMed). For CKKS paramaters, we evaluate different polynomial modulus degree (Poly_mod), coefficient modulus sizes (Coeff_mod), and precision levels. Time(s) for encrypted version shows pre-train/training/total times respectively. Communication costs (Comm_cost) include both pre-training and training rounds. Plain-text FedGCN serves as the baseline for comparison. Communication cost are measured for pre-train communication and training rounds separately.