### v3

# FedGraph: A Research Library and Benchmark for Federated Graph Learning

Abstract
Federated graph learning is an emerging field with significant practical challenges. While algorithms have been proposed to improve the accuracy of training graph neural networks, such as node classification on federated graphs, the system performance is often overlooked, despite it is crucial for real-world deployment. To bridge this gap, we introduce FedGraph, a research library designed for practical distributed training and comprehensive benchmarking of FGL algorithms. FedGraph supports a range of state-of-the-art graph learning methods and includes a monitoring class that evaluates system performance, with a particular focus on communication and computation costs during training. Unlike existing federated learning platforms, FedGraph natively integrates homomorphic encryption to enhance privacy preservation and supports scalable deployment across multiple physical machines with system-level performance evaluation to guide the system design of future algorithms. To enhance efficiency and privacy, we propose a low-rank communication scheme for algorithms like FedGCN that require pre-training communication, accelerating both the pre-training and training phases. Extensive experiments benchmark different FGL algorithms on three major graph learning tasks and demonstrate FedGraph as the first efficient FGL framework to support encrypted low-rank communication and scale to graphs with 100 million nodes [^1] Code available at https://github.com/FedGraph/fedgraph.
1 Introduction
Graph neural networks aim to learn representations of graph-structured data that capture features associated with graph nodes as well as edges between them (Bronstein et al., 2017). Such representations can be used to represent and classify nodes or predict the formation of links in large graphs, or can be used to represent and classify smaller graphs. Most graph applications can modeled as one of three major graph learning problems: node classification (e.g., classifying nodes representing papers in citation networks based on their research topic), link prediction (e.g., recommending the formation of links that represent friendship between users in social networks), or graph classification (e.g. classifying types of proteins in biology, where each protein is represented as a graph). Figure 1 (left) illustrates these three types of graph learning tasks (Benamira et al., 2019; Zhang et al., 2020b).
In practice, graph data is often too large to be trained on a single server or may naturally exist on multiple local clients. For example, graph learning on records of billions of users’ website visits requires significant computational resources, beyond those of a single server. Even if a single server or data center could hold such information, privacy regulations may require that it be stored where it was generated, e.g., the General Data Protection Regulation (GDPR) in Europe and Payment Aggregators and Payment Gateways (PAPG) in India prevent private user data from being shared across international borders. Users may also not want to share their personal data with an external server. In response to these concerns, federated learning has been widely studied as a way to preserve user privacy while training accurate models on data stored at multiple clients (Zhao et al., 2018).
Figure 1: Modeling Applications as Graph Tasks (left) and Federated Graph Training Schematic (right). Node classification aims to predict the labels of grey nodes using structural and feature information from neighboring nodes. Link prediction involves inferring potential future edges between nodes based on existing graph structure and node features. Graph classification assigns labels to entire graphs based on their topology and node information. In federated graph learning, nodes are located across multiple clients with potential cross-client edges. Each client trains a local GNN on its subgraph and exchanges encrypted model updates with a central server, enabling collaborative learning without sharing raw data.
These challenges to training on real-world graphs motivate Federated Graph Learning (FGL) as an important research topic (Liu et al., 2024). The training schematic of such a federated architecture is shown in Figure 1 (right): each client maintains a local graph and a local Graph Neural Network (GNN) model, which can be aggregated with other clients’ local models at a coordinator server. Within this architecture, there are many different FGL algorithms that handle local model updates and aggregation differently, resulting in varying model accuracies and system performance (e.g., runtime, communication costs) (Xie et al., 2021; Zhang et al., 2021; Yao et al., 2024). While evaluating the model accuracy of such algorithms can be easily done with simulations on open graph datasets (Hu et al., 2020), evaluating their real-world system performance requires sophisticated benchmarking platforms that can fairly compare multiple FGL algorithms.
Vanilla FGL benchmark platforms are generally designed for research purposes only and simulate the actions of multiple clients on a single machine (Xie et al., 2021; Li et al., 2024). Thus, they do not capture the real communication and computation costs of training graph models in a federated manner across distinct physical clients. Meanwhile, existing federated learning libraries mainly focus on overcoming client data heterogeneity, with limited support for GNN training. For example, FedScale (Lai et al., 2022) does not support graph models, while FedGraphNN (He et al., 2021) and FederatedScope-GNN (Wang et al., 2022) are not well-maintained and only support basic federated learning methods like FedAvg. Users need to implement the FGL algorithms themselves, including steps unique to FGL tasks such as handling cross-client graph edges that begin and end at different clients. Moreover, none of these platforms natively support FGL enhancements that may be needed in practice, e.g., homomorphic encryption for enhanced privacy (Jin et al., 2023; Yao et al., 2024).
To meet these shortcomings, we introduce FedGraph, a research library to easily train GNNs in federated settings. As shown in Table 1, FedGraph supports various federated training methods of graph neural networks under both simulated and real federated environments, as well as encrypted communication between clients and the central server for model update and information aggregation. We summarize the contributions of FedGraph as follows.
  • FedGraph is the first Python library built for real-world federated graph learning systems, including system optimizations that improve efficiency, scalability, and privacy preservation, as well as multiple state-of-the-art FGL algorithms for easy comparison between them.
  • FedGraph natively supports homomorphic encrypted aggregation to strengthen privacy protection and provides an advanced system-level monitor class that enables detailed analysis of communication and computation costs throughout training.
  • FedGraph proposes a low-rank communication scheme for algorithms like FedGCN to accelerate both pre-train communication and training phases.
  • Extensive experiments benchmark the system performance of various algorithms on three major FGL tasks and showcase the ability to support a privacy-preserving federated system on graphs with 100 million nodes.
Vanilla FGL FedScale FedGraphNN FederatedScope-GNN FedGraph
Distributed Training
Graph Learning
Multiple FGL Algorithms
Cross-Client Edges
Encrypted Aggregation
System Level Profiler
Large Scale ML Optimizations
Table 1: FedGraph Comparison with Existing Frameworks. FedGraph supports distributed training with various FGL algorithms, including algorithms that specifically support cross-client edges and encrypted aggregation. With a system-level profiler, FedGraph can then benchmark existing algorithms and provide optimizations to support training on large (tens of millions of nodes) graphs.
In this paper, we begin with an overview of the system design in Section 2, followed by highlighting the key system components in Section 3. In Section 4, we present a case study demonstrating how FedGraph facilitates the easy design and test of low-rank pre-training communication in FGL. We then benchmark the performance on three major tasks and evaluate its scalability in Sections 5. We finally conclude the paper in Section 6.
2 FedGraph System Design
In this section, we outline the design principles of the FedGraph library and demonstrate how these principles are implemented in our system design (Figure 2).
2.1 Design Principles
The main focus of FedGraph is providing a scalable and privacy-preserving federated graph learning system with ease of use for federated learning researchers and applied scientists in industry. As illustrated in Figure 2 (left), the system architecture is structured according to four design principles.
Optimized usability with simple configurations for simulation and federated training: At the access layer, users configure training settings and can seamlessly switch between local simulation and federated training using the same codebase. With just 10–20 lines of code, users can train a federated GNN model, as we discuss further in Section 2.2. FedGraph abstracts away the complexity of federated graph training, offering a unified platform for training and evaluation.
Benchmark existing federated graph learning methods: In the application layer, FedGraph supports three FGL tasks (node classification, link prediction, and graph classification) and a wide range of state-of-the-art model training algorithms for each task, implemented either by the original authors or library developers. Appendix C lists the currently supported algorithms and datasets.
Extensibility to new datasets and algorithms: In the domain layer, FedGraph separates code into modular components, such as data loaders and training classes. Researchers can extend existing models by inheriting base trainer classes and implementing their own algorithms or data pipelines.
Scalability and privacy preservation for real-world deployment: At the infrastructure layer, FedGraph leverages Ray and Kubernetes to enable distributed computation across edge and cloud environments. Homomorphic encryption can be enabled to secure federated aggregation. This architecture ensures that the system can meet the demands of privacy-preserving, large-scale federated graph learning in realistic settings.
Figure 2: Design Diagram of FedGraph (left) and Quick Start Example (right). The system is organized into four layers: the user access layer, the application layer, the domain layer, and the infrastructure layer. Users only need to focus on the access layer, while the developers can focus on one of the remaining layers based on the domain knowledge.
2.2 FedGraph Use Example
Researchers can easily install the FedGraph library by running pip install fedgraph in Python, configure experiments, and start training federated GNN models. Appendices AB, and D provide more details about the code structure and user API. As illustrated in Figure 2 (right), users can initiate a federated graph learning experiment through a straightforward configuration process. Task, method, and dataset specification: The user specifies the learning task, algorithm, dataset, and client data distribution. FedGraph enforces explicit task-method combinations to ensure compatibility and reproducibility. Training configuration: Key hyperparameters, such as the number of global rounds, local training steps, learning rate, and number of participating trainers, are defined in the configuration. Security and privacy settings: Users can enable homomorphic encryption to ensure privacy-preserving training. Execution: Once settled configuration, calling the API function run_fedgraph(config) launches the experiment. FedGraph then automatically handles data loading, client initialization, and distributed training across clients.
3 FedGraph System Highlights
In this section, we first introduce FedGraph’s monitoring system, which enables usable benchmarking of FGL methods. We then introduce two infrastructure features, FedGraph Homomorphic Encryption for privacy-preserving aggregation and FedGraph Kubernetes for scalable distributed training. Finally, we discuss supported configurations for users to optimize large-scale model training.
3.1 FedGraph Monitoring System
Figure 3: FedGraph Monitoring System Architecture.
The FedGraph Monitoring System, illustrated in Figure 3 (right), uses a Monitor Class to track key system metrics, including running time, CPU/GPU utilization, memory consumption, and communication costs between the server and clients. Examples of the monitoring dashboard during training are provided in Appendix E.
Communication Cost Logging: The system records data transfer rates between the FedGraph User and the API Gateway to assess network performance and identify potential bottlenecks.
Training Time and Accuracy Logging: The monitoring framework captures the training duration and model accuracy throughout the learning process, enabling comparisons across configurations.
CPU/GPU and Memory Usage Tracking: Resource consumption across components, such as the Server Process, Ray Pod AutoScaler, API Gateway, and Object Store, is continuously monitored to provide insights into processing efficiency and to detect memory bottlenecks or potential leaks.
Section 5 further illustrates FedGraph’s resource usage profiles under various experimental settings.
Figure 4: FedGraph Homomorphic Encryption for Secure Aggregation on Encrypted Data.
3.2 FedGraph Homomorphic Encryption
FedGraph supports homomorphic encryption (HE) to enable end-to-end secure computation, ensuring user privacy during training (Figure 4). To our knowledge, it is the first FGL library to incorporate end-to-end HE as a core capability. While prior work has applied HE to specific FGL scenarios (Yao et al., 2024), FedGraph generalizes this support by addressing two key requirements of realistic FGL settings: (i) secure feature aggregation prior to training (Yao et al., 2024), and (ii) secure model aggregation during training (Zhang et al., 2024; Kim et al., 2025).
Pre-Training Aggregation: For algorithms such as FedGCN that require feature aggregation before training, FedGraph applies HE to protect data privacy. Each client encrypts its local node features and transmits the encrypted data to the server. The server performs aggregation directly on the ciphertext and returns the result, which clients then decrypt to obtain the aggregated neighbor features without knowing the original feature.
Federated Aggregation During Training: During model aggregation, FedGraph encrypts each client’s model updates before transmission. The central server then performs homomorphic aggregation on the encrypted updates to compute the global model without accessing any plaintext data (Figure 4(Zhang et al., 2020a; Jin et al., 2023).
Figure 5: Comparison of FedGCN performance on plaintext (blue) and with HE encryption (green).
Homomorphic Encryption Overhead: Figure 5 compares the training time and communication cost for the FedGCN method with and without homomorphic encryption. Since FedGCN involves both pre-training and training-phase aggregation, it serves as a representative case for evaluating the overhead introduced by HE across the full pipeline. As shown in the figure, HE significantly increases communication overhead, particularly during the pre-training phase, where feature matrices are often substantially larger than model parameters. These findings motivate our exploration of communication-efficient techniques in Section 4. Detailed benchmarks, including HE configurations and performance breakdowns, are in Appendix F.
3.3 FedGraph Kubernetes
Figure 6: Kubernetes Service for FedGraph Cluster Management.
To support diverse training workloads and scale to a large number of clients, FedGraph employs Kubernetes as the foundation of its infrastructure layer. Unlike prior platforms (Table 1) that lack native Kubernetes integration, FedGraph is designed to scale efficiently to both large client populations and massive graph datasets. As shown in Figure 6, FedGraph leverages Amazon Web Services Elastic Kubernetes Service to enable flexible and dynamic resource management. A self-managed Kubernetes cluster equipped with GPU instances is deployed to handle computationally intensive tasks. The architecture consists of a master node responsible for orchestration, task scheduling, resource allocation, and cluster health monitoring, while worker nodes execute distributed training processes in parallel. To enhance scalability and cost-efficiency, we integrate the Kubernetes Cluster Autoscaler, which dynamically adjusts worker nodes based on workload demands. This ensures resources are used only when needed, maintaining high efficiency under varying system loads for continuous large-scale execution.
3.4 Optimizations for Scalability
FedGraph is designed to support large-scale federated learning, handling graphs with thousands to millions of nodes through configurable training features.
Client Selection for Large-Scale Federated Learning: FedGraph uses a selective client engagement mechanism, allowing the user to specify a fraction of clients that send the model update at each training round (Li et al., 2019). This approach reduces communication costs at the server and resource usage at the clients, enabling scalability to thousands of clients.
Minibatch Training for Federated Updates: Each client processes only a subset of its local graph, reducing computation and memory demands. This also enables devices with limited resources relative to the size of their training datasets to participate in the training, improving convergence speed without overwhelming network or client resources.
Communication and Resource Optimization: FedGraph optimizes communication, minimizing data transfer between clients and the server. Kubernetes dynamically manages Server Pods and Trainer Pods to meet workload demands, ensuring efficient resource usage as the system scales.
These features collectively enable FedGraph to support federated learning at a massive scale while maintaining resource efficiency and performance.
4 Case Study: Communication and Computation Efficient Federated Node Classification with Low Rank Feature Compression
Efficient communication is a critical challenge in federated graph learning, especially when handling large graphs or deploying privacy-preserving methods such as homomorphic encryption. To address this, we implement low-rank feature compression within FedGraph that significantly reduces communication overhead while preserving model accuracy.
4.1 Architecture Support in FedGraph
FedGraph’s modular design enables the seamless integration of low-rank compression. It separates the pre-training feature aggregation and model training phases, allowing different optimization strategies at each stage. In this case study, we apply HE in both pre-training and model training aggregations, while the low rank compression is applied during pre-training. The HE interface, integrated across multiple algorithms, naturally supports low-rank encrypted aggregation due to its additive structure.
4.2 Low Rank Method for Pre-Train Feature Aggregation
In FedGCN, pre-training communication involves aggregating feature information across clients for nodes with cross-client edges, which may be encrypted and securely aggregated in order to preserve privacy (Yao et al., 2024). We implement a low-rank method for pre-train feature aggregation using client-side projection of feature information to reduce communication cost. The server first generates a random projection matrix $\mathbf{P}\in\mathbb{R}^{d\times k}$ where $k\ll d$ is the specified rank (e.g., $k=100$). The server then distributes $\mathbf{P}$ to all clients. Each client $i$ computes the projected feature matrix $\mathbf{\hat{X}}_{i}=\mathbf{X}_{i}\mathbf{P}$ where $\mathbf{\hat{X}}_{i}\in\mathbb{R}^{n_{i}\times k}$ with $n_{i}$ nodes and then sends $\mathbf{\hat{X}}_{i}$ to the server. The server aggregates these low-rank projections $\mathbf{\hat{X}}_{agg}=\sum_{i=1}^{m}\mathbf{\hat{X}}_{i}$, where $m$ is the number of clients, and distributes the result back to clients. To prevent potential reconstruction of another client’s original features, the projection matrix $\mathbf{P}$ can also be encrypted before being distributed. This adds a layer of protection against inversion attacks on the shared aggregated features. This approach significantly reduces communication overhead in both directions due to only communicating low-rank information. Since our HE interface supports addition on encrypted data, it can safely perform the aggregation procedure on encrypted projected features, preserving client privacy.
4.3 Performance Evaluation
We evaluate the impact of low-rank compression on communication cost, training time, and accuracy in FedGCN using the Cora dataset. Compression ranks range from the full 1433-dimensional features to rank 100, yielding up to 93% reduction. As shown in Figure 7, HE alone introduces significant communication overhead, especially in pre-training, but this is substantially mitigated by applying low-rank projection. Across both plaintext and HE settings, accuracy remains stable even as rank decreases, demonstrating the effectiveness of low-rank compression for communication-efficient and privacy-preserving FGL.
Figure 7: Comparison of communication cost (first and second plots) and training time (third and fourth plots) for FedGCN with different low-rank compression settings. We show results with plaintext and with HE. Each bar represents the total cost divided into pre-training (blue) and training (green) phases, with accuracy plotted as an orange line to show performance trade-offs.
5 Benchmarking FedGraph on Graph Learning Tasks and Scalability
In this section, we benchmark FedGraph’s performance on graph classification, node classification, and link prediction tasks. We then benchmark FedGraph’s performance as we scale the size of the dataset, demonstrating the library’s ability to scale well. Appendix G provides additional results.
5.1 Federated Graph Classification
Figure 8: Accuracy (left), Training Time (middle), and Communication Cost (right) Comparison across Federated Graph Classification Algorithms (10 clients).
The graph classification experiment evaluates the performance of various federated learning algorithms on five benchmark datasets: IMDB-BINARY, IMDB-MULTI, MUTAG, BZR, and COX2. These datasets span diverse graph structures, including social and molecular networks, providing a comprehensive testbed for assessing algorithm effectiveness.
As shown in Figure 8, the evaluation spans 200 training rounds across five datasets. GCFL+ and GCFL+dWs consistently achieve the highest accuracy, particularly on BZR and COX2. However, as illustrated in the middle and right plots, these gains come at the cost of significantly higher training time and communication overhead, especially on complex datasets like IMDB, reflecting their greater computational and communication demands. In contrast, FedAvg offers the lowest communication cost and the shortest, most consistent training time across all datasets, making it a practical choice for latency- and bandwidth-constrained environments.
5.2 Federated Node Classification
Figure 9: Accuracy (left), Training Time (middle), and Communication Cost (right) Comparison across Federated Node Classification Algorithms under $\beta=10000$ (IID data distribution).
We evaluate FedAvg and FedGCN on four node classification benchmarks: Cora, Citeseer, PubMed, and Ogbn-Arxiv. As shown in Figure 9 (left), FedGCN consistently achieves higher accuracy than FedAvg across all datasets. However, this improvement comes at the cost of significantly longer training times, particularly on larger datasets such as Ogbn-Arxiv in Figure 9 (middle).
The communication cost breakdown (Figure 9, right) reveals the source of this overhead: FedGCN incurs substantially higher pre-training communication costs compared to FedAvg. Notably, the observed communication costs closely match the theoretical values, validating the efficiency modeling of these algorithms. This is especially pronounced for Ogbn-Arxiv, where the pre-training phase dominates the total communication volume. While FedAvg has relatively balanced communication between pre-training and training phases, FedGCN’s architecture leads to a disproportionate communication burden during the pre-training aggregation of features across clients. This imbalance helps explain the increased total training time observed for FedGCN on large-scale datasets.
5.3 Federated Link Prediction
Figure 10: AUC (left), Training Time (middle), and Communication Cost (right) Comparison across Federated Link Prediction Algorithms (10 clients).
In the federated link prediction setting, we simulate a scenario where each client holds region-specific data from the Foursquare Global-scale Check-in Dataset (Yang et al., 2016), covering three geographic configurations: US; US and BR; and US, BR, ID, TR, and JP. This setup respects data privacy by avoiding raw data sharing across regions. We evaluate four algorithms, 4D-FED-GNN+, FedLink, STFL, and StaticGNN, across three key metrics: AUC for predictive accuracy, training time for computational efficiency, and communication cost for network efficiency.
As shown in Figure 10, FedLink and STFL achieve the highest AUC scores across all datasets, while StaticGNN and 4D-FED-GNN+ perform moderately well, but show lower AUC on the simpler US-only dataset. In terms of training time, FedLink and STFL incur the highest costs, particularly on the largest regional dataset. In contrast, 4D-FED-GNN+ demonstrates the shortest training times, suggesting suitability for fast iterative training. For communication efficiency, FedLink incurs the highest overhead, especially on large datasets, while StaticGNN consistently exhibits the lowest communication cost, making it the most network-efficient among the evaluated methods.
Figure 11: Accuracy Across Global Training Rounds and CPU Usage for Node Classification Tasks.
5.4 System Performance Monitoring and Resource Utilization
We monitor system performance across 10 training nodes and a central server node using Grafana, which visualizes real-time metrics collected via Prometheus. Figure 11 provides an overview of system behavior during node classification experiments. The first three plots show global test accuracy across training rounds for the Cora, Citeseer, and Pubmed datasets. FedGCN demonstrates significantly faster convergence and higher final accuracy than FedAvg across all datasets. Figure 11 (right) shows CPU usage over time, captured via Grafana. Lighter workloads, such as Cora and Citeseer, result in lower and less frequent CPU utilization spikes. In contrast, Pubmed and Ogbn-Arxiv induce higher and more sustained CPU usage due to their larger graph sizes and increased communication overhead. The pattern of usage spikes aligns with scheduled training rounds, reflecting real-time computational demands at scale.
5.5 Real-World Dataset with Realistic Client Data Distribution
Figure 12: Training Time and Test Accuracy on Ogbn-Papers100M under Different Batch Sizes (800 rounds).
We finally evaluate FedGraph at scale on the Ogbn-Papers100M dataset (over 50GB), one of the largest publicly available graph benchmarks. We use Hugging Face for dataset storage and partitioning, with the Trainer class managing local data loading for each client. The data is distributed across 195 clients  [^2] https://huggingface.co/FedGraph, with node counts assigned following a power law distribution based on country population sizes—mimicking realistic federated environments where larger clients hold more data.
As shown in Figure 12, we assess the effect of varying batch sizes (16, 32, and 64) on training time and test accuracy over 800 training rounds. Training time increases modestly with batch size due to additional per-round computation. Test accuracy remains nearly unchanged, with a slight gain from batch size 16 to 32 and a plateau at 64, suggesting flexibility in batch size selection without compromising model performance. Memory usage remains stable at approximately 17.5 GB, demonstrating FedGraph’s scalability and efficient resource handling in large-scale settings.
6 Conclusion
In this paper, we presented FedGraph, a Python library designed for benchmarking federated graph learning algorithms. Unlike general federated learning platforms, FedGraph supports a diverse set of algorithms and enables systematic comparisons across algorithms, datasets, and system configurations. It features fully distributed training, homomorphic encryption for privacy-preserving scenarios, and a built-in system profiler to measure communication and computation overhead. The modular API allows easy integration of custom datasets and algorithms. Through extensive experiments, including low-rank compression and large-scale training on graphs with up to 100 million nodes, we demonstrate that FedGraph is a practical and scalable tool for real-world FGL evaluation.
While this work focuses on enabling privacy-preserving federated graph learning, future efforts are needed to explore more robust privacy risk assessments, additional optimization strategies, and the inclusion of a broader range of FGL algorithms to expand benchmark coverage and better support industrial deployment.
Appendix A FedGraph Code Structure
Fedgraph library can be separated into six modules:
Data Process Module: data_process.py. This module is responsible for data generation and processing. It should be called before calling the runner, so it is only applicable for node classification and graph classification in the latest version. (In link prediction, the dataset is generated and processed inside the runner.)
Runner Module: federated_methods.py. Task-based runners are defined in this module. For node classification and link prediction, there’s a shared runner that could call inner modules to perform federated graph learning. For graph classification, it will further assign the program to an algorithm-based runner.
Server Classes Module: server_class.py. It defines different task-based server classes including server_NC, server_GC, and server_LP.
Trainer Classes Module: trainer_class.py. It defines different task-based trainer(client) classes including trainer_NC, trainer_GC, and trainer_LP.
Backbone Models Module: gnn_models.py. It defines different backbone model classes. Generally, it is task-based (i.e., each task corresponds to one backbone model), but for some models, their variants are also included and the user could also switch the backbone model or define a new one by itself.
Utility Functions: utils_nc.py, utils_gc.py, utils_lp.py. In the current version, the utility functions for different tasks are located in separate Python modules, which is convenient for developing. In the later versions it might be better to use the shared ‘utils.py‘, and use different markers like “NC”, “GC”, and “LP” to distinguish them.
Appendix B FedGraph API and Runners
When calling the FedGraph API, as shown in Figure 13, Users can simply specify the name of the dataset and algorithm, the API will then call the corresponding data loader class to generate the required data, which is then automatically fed into the appropriate algorithm runner for tasks like Node Classification, Graph Classification, or Link Prediction. Additionally, users can seamlessly add their own datasets or federated graph learning algorithms if needed, as long as they satisfy the form requirements for the specified task.
Figure 13: High-level Demonstration of FedGraph API Design
The function run_fedgraph is a general runner that receives the dataset and the configurations. It will further assign the program to a task-specified runner run_NC , run_GC , or run_LP based on the user’s specification on Task.
{python}
def run_fedgraph(args, data): if args.fedgraph_task == "NC": run_NC(args, data) elif args.fedgraph_task == "GC": run_GC(args, data) elif args.fedgraph_task == "LP": run_LP(args)
Appendix C Supported Algorithms and Datasets
For federated node classification, link prediction, and graph classification, we integrated different datasets and algorithms for each task, shown in Table 2 and Table 3. Researchers can also easily implement new algorithms and add their datasets.
Task Dataset
Node Classification
Cora, Citeseer, Pubmed,
Ogbn-Arxiv, Ogbn-Products, Ogbn-MAG Hu et al. (2020)
Graph Classification
MUTAG, BZR, COX2, DHFR,
PTC-MR, AIDS, NCI1, ENZYMES,
DD, PROTEINS, COLLAB,
IMDB-BINARY, IMDB-MULTI Xie et al. (2021)
Link Prediction FourSquare Yang et al. (2019), WyzeRule Kamani et al. (2024)
Table 2: Supported datasets of node classification, graph classification, and link prediction in federated learning.
Task Algorithm Backbone
Node Classification FedAvg Li et al. (2019) GCN
Distributed GCN GCN
BNS-GCN Wan et al. (2022) GCN
FedSage+ Zhang et al. (2021) GraphSage
FedGCN Yao et al. (2024) GCN,GraphSage
Graph Classification SelfTrain GIN
FedAvg Li et al. (2019) GIN
FedProx Li et al. (2020) GIN
GCFL Xie et al. (2021) GIN
GCFL+ Xie et al. (2021) GIN
GCFL+dWs Xie et al. (2021) GIN
Link Prediction FedAvg Li et al. (2019) GCN
STFL Lou et al. (2021) GCN
FedGNN+ Gürler & Rekik (2022) GCN
FedLink GCN
FedRule Yao et al. (2023) GCN
Table 3: Supported algorithms of node classification, graph classification, and link prediction in federated learning.
Appendix D Runner Workflow
D.1 Runner Workflow for Node Classification Task
For the node classification task, the user specifies the dataset name in the configuration file. The dataset is preprocessed and partitioned across clients based on the chosen federated setting. The data is accessed directly from local storage or via an API, and then a time window may be generated to support temporal learning tasks.
Dataset extraction and preprocessing are managed by the function dataloader_NC in the module data_process.py, which handles both temporal and static graph datasets. Each client holds a subgraph or node features, and training proceeds in a federated manner.
[…]
D.2 Runner Workflow for Graph Classification Task
In the graph classification task, the dataset name can be specified by the user, and it could be either a single dataset or multiple datasets. For single dataset GC, the graphs will be assigned to a designated number of clients; for multiple datasets GC, each dataset will correspond to a client. Dataset generation and preparation are controlled by the function dataloader_GC in the module data_process.py. It will further assign the data process task to the function data_loader_GC_single or data_loader_GC_multiple. All the provided datasets are built-in TUDatasets in the Python library torch_geometric.
[…]
D.3 Runner Workflow for Link Prediction Task
For the link prediction task, we provide a common dataset. The user only needs to specify the country codes. The original datasets are stored in Google Drive, so the user does not need to prepare the dataset by itself. The API will automatically check whether the dataset already exists and download the corresponding one if not.
There are also multiple available algorithms for the link prediction task. However, they share the same set of arguments so that we don’t need to create separate .yaml files. The user could conveniently select different algorithms by directly changing the algorithm field in the configuration file. All the algorithms will share the same runner. The whole workflow is demonstrated in Figure 14(c).
(a) Node classification.
(b) Graph classification.
(c) Link prediction.
Figure 14: Federated graph training workflows for (a) node classification, (b) graph classification, and (c) link prediction tasks.
Appendix E System Measurement Dashboard
In Figure 15, the FedGraph project monitors system performance by integrating Prometheus and Grafana within a Kubernetes cluster. Prometheus collects metrics from Ray clusters, while Grafana visualizes these metrics through embedded panels, allowing real-time monitoring and analysis.
Figure 15: FedGraph Monitoring System.
Appendix F Benchmark Homomorphic Encryption
We provide a comprehensive guide for the configuration of Homomorphic Encryption in FedGraph to provide security guarantees corresponding to the graph structure and size.
F.1 Parameter Configuration
In Table 4, we present the key parameters for configuring CKKS homomorphic encryption. The selection of the parameters is based on the dataset size and desired security level. Different parameter combinations create tradeoffs between computational overhead, communication cost, and precision. We present the default setting and parameters selection range to guide the user in selecting an appropriate combination that achieves the balance between objectives.
Parameter Default Value Description Range
scheme CKKS Encryption scheme type N/A
polynomial modulus degree 16384 Maximum degree of polynomials used to represent encrypted data ($N\geq 2\times$ max(nodes, features)) 4096, 8192, 16384, 32768
coefficient modulus bit size [60, 40, 40, 40, 60] Bit size for coefficient modulus that controls precision Array of integers in the range $[20,60]$
global scale $2^{40}$ Global scale factor for encoding precision $2^{30}$, $2^{40}$, $2^{50}$, etc.
security level 128 Bit security level 128, 192, 256
Table 4: TenSEAL Homomorphic Encryption Configuration Parameters. This table shows the key parameters for configuring the CKKS encryption scheme in the FedGraph library, including their default values, descriptions, and available ranges.
F.2 Microbenchmark
We then provide the microbenchmark of HE on federated graph training in Table 5. The experiments are conducted on a 2-layer FedGCN for node classification tasks, running 100 global rounds with default settings in Cora. For CKKS parameters, we evaluate different polynomial modulus degrees (Poly_mod), coefficient modulus sizes (Coeff_mod), and precision levels. Time(s) for the encrypted version shows pre-train/training/total times respectively. Communication costs (Comm_cost) include both pre-training and training rounds.
Dynamic Precision: We adjust encryption parameters based on graph sizes and the numerical precision needed. For graphs like Cora, a polynomial modulus degree of 16384 with precision $2^{40}$ satisfies the ideal accuracy, while increased value provides more precise security protection.
Communication Cost Optimization: We employ several strategies to manage the communication overhead inherent in HE operation. The selection of coefficient modulus chain, [60,40,40,40,60], etc., enables efficient depth management for multiple HE operations. Depending on specific dataset characteristics (sparse matrix, larger datasets, etc.), we also employ efficient encryption methods to optimize communication cost and balance the performance.
When comparing among datasets, we observe that HE maintains equivalent accuracy across different parameter selections, as long as they satisfy the modulus requirement. If a smaller than required parameter size is used, the accuracy drops sharply which indicates not valid encryption.
Method Poly _mod Coeff_mod Precision Dataset Time(s) Comm_cost (MB) Accuracy
FedGCN (plaintext) N/A N/A N/A Cora 13.29 59.21 0.783 $\pm$ 0.07
FedGCN (HE) 16384 [60,40,40,40,60] $2^{40}$ Cora 27.71/23.22/56.05 3279.15 0.779 $\pm$ 0.08
FedGCN (HE) 32768 [60,40,40,40,60] $2^{50}$ Cora 29.44/36.17/71.76 4434.58 0.781 $\pm$ 0.08
FedGCN (plaintext) N/A N/A N/A Citeseer 20.39 187.99 0.658 $\pm$ 0.06
FedGCN (HE) 8192 [60,40,40,60] $2^{40}$ Citeseer 79.35/30.63/113.08 5791.42 0.660 $\pm$ 0.07
FedGCN (HE) 16384 [60,40,40,40,60] $2^{40}$ Citeseer 123.22/45.8/173.40 8084.50 0.652 $\pm$ 0.06
FedGCN (plaintext) N/A N/A N/A PubMed 25.78 150.43 0.774$\pm$ 0.12
FedGCN (HE) 8192 [60,40,40,60] $2^{40}$ PubMed 70.28/19.73/93.18 3612.60 0.757 $\pm$ 0.19
FedGCN (HE) 16384 [60,40,40,40,60] $2^{40}$ PubMed 123.22/45.8/173.40 8084.50 0.769 $\pm$ 0.13
Table 5: Microbenchmark of FedGCN under Homomorphic Encryption with different CKKS scheme parameters. The experiments are conducted on a 2-layer FedGCN for node classification tasks, running 100 global rounds with default settings across three datasets (Cora, Citeseer, PubMed). For CKKS paramaters, we evaluate different polynomial modulus degree (Poly_mod), coefficient modulus sizes (Coeff_mod), and precision levels. Time(s) for encrypted version shows pre-train/training/total times respectively. Communication costs (Comm_cost) include both pre-training and training rounds. Plain-text FedGCN serves as the baseline for comparison. Communication cost are measured for pre-train communication and training rounds separately.
Appendix G Additional Experiments
G.1 Mini-Batch for Large Dataset
FedGraph inherently supports a mini-batch graph splitter, enabling efficient graph-level processing. Our experiments on the ogbn-arxiv dataset examined the impact of varying batch sizes (16, 32, 64) on test accuracy, training time, memory usage, and communication cost across 10 clients in a federated setting.
Results in Figure 16 show test accuracy peaks at a batch size of 64, with improvements leveling off beyond 32. Training time and memory demands increase with batch size, though pretrain memory remains stable.
Figure 16: Pretrain, Train time, Communication Cost, Accuracy of FedGCN on Ogbn-Arxiv in Different Batch Sizes (200 rounds)
G.2 Increasing the Number of Clients
Figure 17: Training Time, Communication Cost, Test Accuracy on Ogbn-Arxiv in Large Number of Clients. All experiments run on 10 AWS instances. 1000 trainers take long since it runs sequentially on 10 instances.
To better test scalability and fit real-world data, we increase the number of clients to 1000. In Figure 17, we observe that as the number of clients increases, the overall training time grows significantly due to sequential running on 10 instances, added communication overhead and the need for increased synchronization among clients.
As we scale from 10 to 1000 clients under a fixed IID Beta value, there is a small decline in accuracy, likely due to the increased data heterogeneity each client possesses. The communication cost also escalates notably with more clients, highlighting the trade-off between parallelism and efficiency in federated settings. This experiment shows the system’s ability to handle large-scale client distributions while revealing the resources required to maintain accuracy and efficiency.